|x^2 - 4x - 4 | = 8
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39633)
Answer by ikleyn(52946)

.
|x^2 - 4x - 4 | = 8    (1)


The equation (1) is equivalent to this:


    Find the solution set to this set of equations:

        EITHER   x^2 - 4x - 4 = 8       (2)
        OR       x^2 - 4x - 4 = -8.     (3)


Notice that these two equations, (2) and (3), are connected by the logical term "OR",

which means that the set of solutions to equation (1)  is the UNION  of solutions to equations  (2)  and (3).


Now,

    x^2 - 4x - 4 = 8   ====>  x^2 - 4x - 12 = 0  ====>  (x+2)\*(c-6) = 0  ====>  the solutions are  x = -2  and/or  x = 6,

and

    x^2 - 4x - 4 = -8  ====>  x^2 - 4x + 4 = 0  ====>  (x-2)^2 = 0  ====>  the solution is  x = 2.


Thus the solution to the given equation (1) is the set  {-2, 2, 6}.




Plot y = |x^2 - 4x - 4| (red) and y = 8 (green)


--------------
On solving  Absolute Value equations  there is a bunch of lessons in this site
    - Absolute Value equations
    - HOW TO solve equations containing Linear Terms under the Absolute Value sign. Lesson 1
    - HOW TO solve equations containing Linear Terms under the Absolute Value sign. Lesson 2
    - HOW TO solve equations containing Linear Terms under the Absolute Value sign. Lesson 3
    - HOW TO solve equations containing Quadratic Terms under the Absolute Value sign. Lesson 1
    - HOW TO solve equations containing Quadratic Terms under the Absolute Value sign. Lesson 2
    - OVERVIEW of lessons on Absolute Value equations

Read them attentively and become an expert in this area.


Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
"Solving Absolute values equations".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.